john and charles are linemen. John can string 10 miles of line in 3 days, while together they can string it in 1 day. how long would it take charles alone to string the same line?
?

Answer by ptaylor(2048)

Let x=amount of time it takes Charles alone to string the same line
Then Charles works at the rate of 1/x of the line per day
John can string 1/3 of the line per day
Together John and Charles can string 1/x + 1/3 of the line per day
But we are told that together they can do the job in 1 day, so:
1/x + 1/3 =1 multiply each term by 3x
3+x=3x
2x=3
x=1.5 days
CK
1/(3/2) +1/3=1
2/3+1/3=1
1=1
Hope this helps---ptaylor